(11 %) Problem 8: Suppose you have a 9.00 V battery, a 2.8 μF capacitor, and a 8.45 μF capacitor. 25% Part (a) Find the total charge stored in the system if the capacitors are connected to the battery in series in C. sin cosO tan HOME acosO atan)acotan)sinh) 0 coshO tanh) cotanhO DegreesRadians END BACKSPACE DEL CLEAR Submit Hint I give up! Hints: 1 % deduction per hint. Hints remaining: 2 Feedback: 1% deduction per feedback 25% Part (b) Find the energy stored in the system if the capacitors are connected to the battery in series in J. 25% Part (c) Find the charge if the capacitors are connected to the battery in parallel in C. 25% Part (d) Find the energy stored if the capacitors are connected to the battery in parallel in J.
Answer #1

Fos seuies lonnected Capacito , eauivalent capacitance e9 19-934c 2. Etwvnlent CA f ad ten re pa.ai tel tern neded afhz刎
Answer #2

a) for the capacitors in series

1/C = 1/2.8 + 1/8.45

C = 2.103 uF

total charge stored = Ceq * V

total charge stored = 2.103 * 9

total charge stored = 18.93 uC

total charge stored = 1.893 *10-5 C

b)

total energy stored = 0.50 C * V^2

total energy stored = 0.50 * 2.103 *10^-6 * 9^2

total energy stored = 8.52 *10-5 J

c) in parallel

total charge = (2.8 + 8.45) *10^-6 * 9

total charge = 1.01 *10-4 C

d) total energy stored = 0.50 * 2.8 *10^-6 * 9^2 + 0.50 * 8.45 *10^-6 * 9^2

total energy stored = 4.55 * 10-4 J